OpenStudy (loser66):

evaluate int_(0--> 2) int_(0--> sqrt(y)) x e^y dxdy step by step, please

OpenStudy (anonymous):

need help

terenzreignz (terenzreignz):

hmm... \[\huge \int\limits_{0}^2\int\limits_{0}^\sqrt yxe^y \ dxdy\]

terenzreignz (terenzreignz):

Okay, so, first thing you do is integrate this part...\[\huge \int\limits_{0}^2\color{red}{\int\limits_{0}^\sqrt yxe^y \ dx}dy\]pretending y is just a constant.

OpenStudy (loser66):

= int. x^2/2 * e^y dy

terenzreignz (terenzreignz):

Hold on there... \[\huge \int\limits_{0}^{\sqrt y}xe^y dx=\left.\frac{x^2e^y}{2}\right]_0^{\sqrt y} \]

OpenStudy (loser66):

int. x^2/2*e^y ( replace by limits) dy ,sorry

OpenStudy (loser66):

yes

terenzreignz (terenzreignz):

Okay, so can you evaluate this part first...\[\huge \int\limits_{0}^{\sqrt y}xe^y dx=\color{blue}{\left.\frac{x^2e^y}{2}\right]_0^{\sqrt y}}\]

OpenStudy (anonymous):

wate im lost now

OpenStudy (loser66):

= (y e^y)/2